Hello-Today on test, I was unable to prove that x=2. Given: 1/2x + 1 = 2x -2.
I know x=2 but kept coming up with x=1 when I did all of the steps. I must be doing something wrong. Thanks.

Answer by nerdybill(7384)

If you had:
(1/2)x + 1 = 2x -2
and you wanted to prove that x=2. Simply, plug it the value:
.
(1/2)x + 1 = 2x -2
(1/2)(2) + 1 = 2(2) -2
(2/2) + 1 = 2(2) -2
1 + 1 = 4 -2
2 = 2 (checks)